What is the answer to: 3/5×2/3+1/4×1/3
1
Expert's answer
2018-01-23T15:12:08-0500
3/5∙2/3+1/4∙1/3=2/5+1/12=(2∙12)/(5∙12)+(1∙5)/(5∙12)=(24+5)/60=29/60.